a potential difference of 10v causes a current of 5a to flow through a lamp. if the bulb is switched on for 20 seconds, how much energy is transferred ?


1
Expert's answer
2021-02-10T12:52:35-0500



 Transferred energy = "I*V*t"


I= 5A

V=5 v


t=20 sec

 Transferred energy = 5 * 5 * 20 = 25 * 20 = 500 J

Answer : 500 J